Two charges, magnitude + 2 x 10-6 C each are 60cm apart. Find the magnitude of the force exerted by these charges on a third charge of magnitude + 4 x 10-6 C that is 50 cm away from each of the first two charges.
1
Expert's answer
2020-03-13T10:53:07-0400

Solution. Since charges of the same sign, two charges repel the third with the forces F1 and F2 equal in absolute value



According to the conditions of the problem r=0.6m; r1=r2=0.5m; q1=q2=2x10-6С; q3=4x10-6C.



"F_1=F_2=\frac{kq_1q_3}{r_1^2}=\frac{kq_2q_3}{r_2^2}=\frac{9\times10^9\times 2\times 10^{-6}\times 4\times 10^{-6}}{0.5^2}=0.288N"

The angle between the vectors is equal to the angle ACB. Therefore, the resulting force is


"\vec{F}=\vec{F_1}+\vec{F_2}"

Using the cosine theorem the magnitude of the resulting force


"F^2=F_1^2+F_2^2-2F_1F_2cos(180^0-\u2220ACB)""F=\sqrt{F_1^2+F_2^2-2F_1F_2cos(180^0-∠ACB)}=\sqrt{2F_1^2+2F_1^2cos∠ACB}"

"F=F_1\sqrt{2+2cos∠ACB}"

Using the cosine theorem for ∆ABC get


"r^2=r_1^2+r_2^2-2r_1r_2cos\u2220ACB""cos∠ACB=\frac{r_1^2+r_2^2-r^2}{2r_1r_2}=\frac{0.5^2+0.5^2-0.6^2}{2\times0.5\times0.5}=0.28"

As result


"F=0.288\sqrt{2+2\times 0.28}=0.4608N"

Answer. 0.4608N